# Fast Online $\ell_{0}$ Elastic Net Subspace Clustering via A Novel Dictionary Update Strategy

### v1

@@ Abstract
With the rapid growth of data volume and the increasing demand for real-time analysis, online subspace clustering has emerged as an effective tool for processing dynamic data streams. However, existing online subspace clustering methods often struggle to capture the complex and evolving distribution of such data due to their reliance on rigid dictionary learning mechanisms. In this paper, we propose a novel $\ell_{0}$ elastic net subspace clustering model by integrating the $\ell_{0}$ norm and the Frobenius norm, which owns the desirable block diagonal property. To address the challenges posed by the evolving data distributions in online data, we design a fast online alternating direction method of multipliers with an innovative dictionary update strategy based on support points, which are a set of data points to capture the underlying distribution of the data. By selectively updating dictionary atoms according to the support points, the proposed method can dynamically adapt to the evolving data characteristics, thereby enhancing both adaptability and computational efficiency. Moreover, we rigorously prove the convergence of the algorithm. Finally, extensive numerical experiments demonstrate that the proposed method improves clustering performance and computational efficiency, making it well-suited for real-time and large-scale data processing tasks.
@@
$\ell_{0}$ elastic net , online algorithm , subspace clustering , support point.
@@ 2.2 Optimality Conditions
For convenience, denote $h(X):=\frac{1}{2}\lVert Z-DX\rVert_{F}^{2}+\frac{\lambda_{2}}{2}\lVert X\rVert_{F}^{2}$. Then, $\nabla h(X)=(D^{T}D+\lambda_{2}I)X-D^{T}Z$. It is easy to verify that $h$ is strongly convex and $r$-smooth with $r=\lVert D^{T}D+\lambda_{1}I\rVert_{2}$. For a $r$-smooth function $h$, the following inequality holds
@@
where $U=[\mathbf{u}_{1},\ldots,\mathbf{u}_{n}]\in\mathbb{R}^{m\times n}$ is the Lagrangian multiplier. We say that $(X^{*},Y^{*};U^{*})$ is a KKT point of model (3) if it satisfies
@@
where $Prox_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(\cdot\right)$ is the proximal operator of $\alpha\lambda_{1}\|\cdot\|_{0}$.
It is worth mentioning that $Prox_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(\cdot\right)$ is a multi-valued operator due to the nonconvexity of $\alpha\lambda_{1}\|\cdot\|_{0}$. According to the definition of the proximal operator, for a given matrix $A$, we know that, for any $i\in\{1,2,\ldots,n\}$,
$\left[Prox_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(A\right)\right]_{ij}=\arg\min_{X_{ij}}\frac{1}{2}\|X_{ij}-A_{ij}\|_{2}^{2}+\alpha\lambda_{1}\mathbb{I}(A_{ij}\neq 0).$ (8)
@@
$\left[Prox_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(W\right)\right]_{ij}=\begin{cases}0,&\frac{1}{2}A_{ij}^{2}<\alpha\lambda_{1},\\ \{W_{ij},0\},&\frac{1}{2}A_{ij}^{2}=\alpha\lambda_{1},\\ W_{ij},&\frac{1}{2}A_{ij}^{2}>\alpha\lambda_{1}.\\ \end{cases}$ (9)
@@
    $\left|Y^{*}_{ij}\right|\geq\sqrt{2\alpha\lambda_{1}}$ or $Y^{*}_{ij}=0,\forall i,j$;
@@
    $\left|X^{*}_{ij}\right|\geq\sqrt{2\alpha\lambda_{1}}$ or $X^{*}_{ij}=0,\forall i,j$.
@@
$\alpha_{1}^{*}:=\frac{\epsilon_{*}}{2\lambda_{1}\lVert Y^{*}\rVert_{0}+2\sqrt{\lambda_{1}^{2}\lVert Y^{*}\rVert_{0}^{2}+\epsilon_{*}\lVert\nabla h(X^{*})\rVert_{F}^{2}}},\alpha_{2}^{*}:=\frac{1}{\sqrt{n\|D^{T}D+\lambda_{2}I\|_{F}^{2}}}.$
@@ 3.1 Online Optimization Algorithm
$\mathbf{y}_{i}^{k+1}=\arg\min_{\mathbf{y}_{i}}\mathcal{L}_{\sigma}(\mathbf{y}_{i},\mathbf{x}_{i},\mathbf{u}_{i}),$ (12)
$\mathbf{x}_{i}^{k+1}=\arg\min_{\mathbf{x}_{i}}\mathcal{L}_{\sigma}(\mathbf{y}_{i},\mathbf{x}_{i},\mathbf{u}_{i}),$ (13)
@@
$\mathbf{x}_{i}^{k+1}=\left[D^{T}D+(\lambda_{2}+\sigma)I\right]^{-1}\left(D^{T}\mathbf{z}_{i}-\mathbf{u}_{i}^{k}+\sigma\mathbf{y}_{i}^{k+1}\right).$ (18)
@@ 5             Update $\mathbf{x}_{i}^{k+1}$ by (18);
Algorithm 1 Online ADMM for Sloving (3)
@@
where $\kappa=\frac{(2+\sigma)(\gamma+\lambda_{2})+\sigma^{2}}{2\sigma}$ and $\gamma=\lambda_{min}(D^{T}D)$.
@@
Let $\sigma\geq 2\|D^{T}D+\lambda_{1}I\|_{2}$ and $\{\mathbf{w}_{i}^{k}\}$ be the sequence generated by Algorithm 1. Then, $\{\mathbf{w}_{i}^{k}\}$ is bounded. Moreover, the following statement holds
@@
Let $\sigma\geq 2\|D^{T}D+\lambda_{1}I\|_{2}$ and $\{\mathbf{w}_{i}^{k}\}$ be the sequence generated by Algorithm 1. Then any accumulation point of $\{\mathbf{w}_{i}^{k}\}$ is a P-stationary point and also a uniquely local minimizer of (3).
Next, we analyze the complexity of Algorithm 1. Since Algorithm 1 is an online method that processes samples in a streaming manner, its computational complexity per sample is given by $\mathcal{O}\left(m(p^{2}+m^{2})\right)$, where $m$ and $p$ are the numbers of dictionary samples and features respectively. This complexity arises from the following three main components. Firstly, updating $\mathbf{y}_{i}$ primarily involves element-wise addition and a hard threshold operation, resulting in a computational complexity of $\mathcal{O}\left(m\right)$. Secondly, updating $\mathbf{x}_{i}$ is the most computationally intensive step. This step involves solving an equation that requires matrix operations such as calculating $D^{T}D$, adding a scaled identity matrix, and computing the inverse of the resulting matrix. Hence, the overall complexity for updating $\mathbf{x}_{i}$ is $\mathcal{O}\left(m(p^{2}+m^{2})\right)$. Finally, updating $\mathbf{u}_{i}$ only involves simple vector addition, which has a complexity of $\mathcal{O}\left(m\right)$.
3.2 Dictionary Update Based on Support Point
@@
We begin by recalling the definition of the support point proposed in [30].
@@
For a fixed-point set $Z\in\mathbb{R}^{p\times n}$, the support points of $Z$ are defined as
$\{\boldsymbol{\xi}_{i}\}_{i=1}^{m}=\arg\min_{\mathbf{d}_{1},\ldots,\mathbf{d}_{m}}\frac{2}{mn}\sum_{i=1}^{m}\sum_{l=1}^{n}\|\mathbf{z}_{l}-\mathbf{d}_{i}\|_{2}-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\|\mathbf{d}_{i}-\mathbf{d}_{j}\|_{2}.$ (21)
@@
Assume the dataset $Z\in\mathbb{R}^{p\times n}$ is drawn from $c$ independent subspaces $\{\mathcal{S}_{i}\}_{i=1}^{c}$ with the corresponding dimension $\{d_{i}=dim(\mathcal{S}_{i})\}_{i=1}^{c}$. Then, it needs at least $\sum_{i=1}^{c}\left(d_{i}+1\right)$ points to adequately capture the distribution of the entire data.
@@
$\min_{\mathbf{d}_{1},\ldots,\mathbf{d}_{m}}\frac{2}{mn}\sum_{i=1}^{m}\sum_{l=1}^{n}\|\mathbf{z}_{l}-\mathbf{d}_{i}\|_{2}-\frac{1}{m^{2}}\left[\sum_{i=1}^{m}\sum_{j=1}^{m}\left(\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}+\frac{2(\mathbf{d}_{i}-\mathbf{d}^{{}^{\prime}}_{j})^{T}(\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j})}{\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}}\right)\right].$ (22)
@@
$\begin{split}\min_{\mathbf{d}_{1},\ldots,\mathbf{d}_{m}}\frac{2}{mn}\sum_{i=1}^{m}&\sum_{l=1}^{n}\left[\frac{\|\mathbf{z}_{l}-\mathbf{d}_{i}\|_{2}^{2}}{2\|\mathbf{z}_{l}-\mathbf{d}^{{}^{\prime}}_{i}\|_{2}}+\frac{\|\mathbf{z}_{l}-\mathbf{d}^{{}^{\prime}}_{i}\|_{2}}{2}\right]\\ &-\frac{1}{m^{2}}\left[\sum_{i=1}^{m}\sum_{j=1}^{m}\left(\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}+\frac{2(\mathbf{d}_{i}-\mathbf{d}^{{}^{\prime}}_{j})^{T}(\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j})}{\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}}\right)\right].\end{split}$ (23)
@@
Assume $\mathcal{D}$ is compact and convex and $\{D^{k}\}$ is the sequence generated by Algorithm 2. If $D_{0}\subseteq\mathcal{D}$, any limit point of $\{D^{k}\}$ is a local minimizer of (21).
@@ 3.3 Online Algorithm with Dictionary Update
Algorithm 3 ONESC Algorithm with Support Points
@@ 7                   Update $D$ by Algorithm 2 with $\hat{Z}$;
Algorithm 4 ONESC Algorithm with Dictionary Update (OENSC-S)
@@ Acknowledgment
According to [17], we only need to verify that $g(X)=\lambda_{1}\lVert X\rVert_{0}+\frac{\lambda_{2}}{2}\lVert X\rVert_{F}^{2}$ satisfies the following two conditions: (1) $g(P^{T}XT)=g(X)$ for any permutation matrix $P$; (2) $g(X)\geq g(X^{B})$ with $X=\left(\begin{matrix}X_{1}&X_{2}\\ X_{3}&X_{4}\end{matrix}\right)$ and $X=\left(\begin{matrix}X_{1}&0\\ 0&X_{4}\end{matrix}\right)$, in which the equality holds if and only if $X=X^{B}$.
@@
$g(P^{T}XP)$ $=\lambda_{1}\lVert P^{T}XP\rVert_{0}+\frac{\lambda_{2}}{2}\lVert P^{T}XP\rVert_{F}^{2}$
$=\lambda_{1}\sum_{i,j}\mathbb{I}\left[(P^{T}XP)_{ij}\neq 0\right]+\frac{\lambda_{2}}{2}\sum_{i,j}\left[\left(P^{T}XP\right)_{i,j}\right]^{2}$
$=\lambda_{1}\sum_{i,j}\mathbb{I}\left[P^{T}(XP)_{ij}\neq 0\right]+\frac{\lambda_{2}}{2}\sum_{i,j}\left[P^{T}\left(XP\right)_{i,j}\right]^{2}$
$=\lambda_{1}\sum_{i,j}\mathbb{I}\left[P^{T}(X)_{ij}\neq 0\right]+\frac{\lambda_{2}}{2}\sum_{i,j}\left[P^{T}\left(X\right)_{i,j}\right]^{2}$
@@
It follows from (9) that the second equation of (7) can be written as
$Y^{*}_{ij}=\left[Prox_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(Y^{*}+\alpha U^{*}\right)\right]_{ij}=\begin{cases}0,&(Y^{*}+\alpha U^{*})_{ij}\leq\sqrt{2\alpha\lambda_{1}},\\ (Y^{*}+\alpha U^{*})_{ij},&(Y^{*}+\alpha U^{*})_{ij}>\sqrt{2\alpha\lambda_{1}}.\end{cases}$
@@
$0=Y^{*}-Prox_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(Y^{*}+\alpha U^{*}\right)=\left(\begin{matrix}Y^{*}_{T}\\ Y^{*}_{\bar{T}}\end{matrix}\right)-\left(\begin{matrix}0\\ Y^{*}_{\bar{T}}+\alpha U^{*}_{\bar{T}}\end{matrix}\right)=\left(\begin{matrix}Y^{*}_{T}\\ \alpha U^{*}_{\bar{T}}\end{matrix}\right),$
where $T=\{(i,j):Y^{*}_{ij}\leq\sqrt{2\alpha\lambda_{1}}\}$ and $\bar{T}$ is the complement set of $T$. Thus, $Y^{*}_{ij}=0$ for any $(i,j)\in T$ and $\left|Y^{*}_{ij}\right|\geq\sqrt{2\alpha\lambda_{1}}$ for any $(i,j)\in\bar{T}$, which implies the assertion (a). Then the assertion (b) holds by the third equation of (7). This completes the proof.
@@
By the definition of the proximal operator, it follows from the second equation in (7) that $0\in\alpha\lambda_{1}\partial\lVert Y^{*}\rVert_{0}+Y^{*}-(Y^{*}+\alpha U^{*}).$ This together with $\alpha>0$ immediately yields that $(Y^{*},X^{*},U^{*})$ is a KKT point. Next, we demonstrate that it is also a local optimal solution. Since $h$ is a strongly convex function, [44, Theorem 3.2] leads to it is also a uniquely local minimizer. The proof is completed.
@@
Let $\gamma=\lambda_{min}(D^{T}D)$, it derives
$\mathcal{L}_{\sigma}(\mathbf{y}_{i}^{k+1},\mathbf{x}_{i}^{k+1},\mathbf{u}_{i}^{k+1})-\mathcal{L}_{\sigma}(\mathbf{y}_{i}^{k+1},\mathbf{x}_{i}^{k+1},\mathbf{u}_{i}^{k})$ $\leq-\frac{1}{\sigma}\|\left(D^{T}D+\lambda_{2}I\right)\left(\mathbf{x}_{i}^{k+1}-\mathbf{x}_{i}^{k}\right)\|_{2}^{2}$
@@
Secondly, we show that any limit points of $\{\mathbf{w}_{i}\}$ is a P-stationary point of problem (3). By taking the limit on both sides of the update formula for the Lagrangian multiplier, it can be reached that $\mathbf{x}_{i}^{*}-\mathbf{y}_{i}^{*}=0$, which demonstrates the feasibility of $\mathbf{w}_{i}^{*}$. Then, the first equation in (7) can be obtained directly from the $\mathbf{x}_{i}$-subproblem. Furthermore, $\mathbf{y}_{i}^{k_{j}}=Prox_{\frac{\lambda_{1}}{\sigma}\|\cdot\|_{0}}(\mathbf{x}_{i}^{k_{j}}+\mathbf{u}_{i}^{k_{j}}/\sigma)$ can be derived by taking limits on both sides of (16). Thus, $\{\mathbf{w}_{i}^{*}\}$ is a stationary point of problem (3).

### v2

@@ Abstract
With the rapid growth of data volume and the increasing demand for real-time analysis, online subspace clustering has emerged as an effective tool for processing dynamic data streams. However, existing online subspace clustering methods often struggle to capture the complex and evolving distribution of such data due to their reliance on rigid dictionary learning mechanisms. In this paper, we propose a novel $\ell_{0}$ elastic net subspace clustering model by integrating the $\ell_{0}$ norm and the Frobenius norm, which owns the desirable block diagonal property. To address the challenges posed by the evolving data distributions in online data, we design a fast online alternating direction method of multipliers with an innovative dictionary update strategy based on support points, which are a set of data points to capture the underlying distribution of the data. By selectively updating dictionary atoms according to the support points, the proposed method can dynamically adapt to the evolving data characteristics. Moreover, we rigorously prove the convergence of the algorithm. Finally, extensive numerical experiments demonstrate that the proposed method improves clustering performance and computational efficiency, making it well-suited for real-time and large-scale data processing tasks.
@@
$\ell_{0}$ elastic net , online algorithm , subspace clustering , support points.
@@ 2.2 Optimality Conditions
For convenience, denote $h(X):=\frac{1}{2}\lVert Z-DX\rVert_{F}^{2}+\frac{\lambda_{2}}{2}\lVert X\rVert_{F}^{2}$. Then, $\nabla h(X)=(D^{\top}D+\lambda_{2}I)X-D^{\top}Z$. It is easy to verify that $h$ is strongly convex and $r$-smooth with $r=\lVert D^{\top}D\rVert_{2}+\lambda_{2}$. For a $r$-smooth function $h$, the following inequality holds
@@
where $U=[\mathbf{u}_{1},\ldots,\mathbf{u}_{n}]\in\mathbb{R}^{m\times n}$ is the Lagrangian multiplier. We say that $(X^{*},Y^{*},U^{*})$ is a KKT point of model (3) if it satisfies
@@
where $\operatorname{Prox}_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(\cdot\right)$ is the proximal operator of $\alpha\lambda_{1}\|\cdot\|_{0}$.
It is worth mentioning that $\operatorname{Prox}_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(\cdot\right)$ is a multi-valued operator due to the nonconvexity of $\alpha\lambda_{1}\|\cdot\|_{0}$. According to the definition of the proximal operator, for a given matrix $A$, we know that, for any $i\in\{1,2,\ldots,n\}$,
$\left[\operatorname{Prox}_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(A\right)\right]_{ij}=\underset{X_{ij}}{\arg\min}\frac{1}{2}\|X_{ij}-A_{ij}\|_{2}^{2}+\alpha\lambda_{1}\mathbb{I}(X_{ij}\neq 0).$ (8)
@@
$\left[\operatorname{Prox}_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(A\right)\right]_{ij}=\begin{cases}0,&\frac{1}{2}A_{ij}^{2}\leq\alpha\lambda_{1},\\ A_{ij},&\frac{1}{2}A_{ij}^{2}>\alpha\lambda_{1}.\\ \end{cases}$ (9)
@@
    $\left|Y^{*}_{ij}\right|>\sqrt{2\alpha\lambda_{1}}$ or $Y^{*}_{ij}=0,\forall i,j$;
@@
    $\left|X^{*}_{ij}\right|>\sqrt{2\alpha\lambda_{1}}$ or $X^{*}_{ij}=0,\forall i,j$.
@@
$\alpha_{1}^{*}:=\frac{\epsilon_{*}}{2\lambda_{1}\lVert Y^{*}\rVert_{0}+2\sqrt{\lambda_{1}^{2}\lVert Y^{*}\rVert_{0}^{2}+\epsilon_{*}\lVert\nabla h(X^{*})\rVert_{F}^{2}}},\alpha_{2}^{*}:=\frac{1}{\sqrt{n}\|D^{\top}D+\lambda_{2}I\|_{F}}.$
@@ 3.1 Online Optimization Algorithm
$\mathbf{y}_{i}^{k+1}=\underset{\mathbf{y}_{i}}{\arg\min}\mathcal{L}_{\sigma}(\mathbf{y}_{i},\mathbf{x}_{i},\mathbf{u}_{i}),$ (12)
$\mathbf{x}_{i}^{k+1}=\underset{\mathbf{x}_{i}}{\arg\min}\mathcal{L}_{\sigma}(\mathbf{y}_{i},\mathbf{x}_{i},\mathbf{u}_{i}),$ (13)
@@
$\mathbf{x}_{i}^{k+1}=\left[D^{\top}D+(\lambda_{2}+\sigma)I\right]^{-1}\left(D^{\top}\mathbf{z}_{i}-\mathbf{u}_{i}^{k}+\sigma\mathbf{y}_{i}^{k+1}\right).$ (18)
@@ 5             Update $\mathbf{x}_{i}^{k+1}$ by (18);
Algorithm 1 Online ADMM for Solving (3)
@@
where $\kappa=\frac{(2+\sigma)(\gamma+\lambda_{2})+\sigma^{2}}{2\sigma}$ and $\gamma=\lambda_{min}(D^{\top}D)$.
@@
Let $\{\mathbf{w}_{i}^{k}\}$ be the sequence generated by Algorithm 1 and $\sigma\geq r$. Then, $\{\mathbf{w}_{i}^{k}\}$ is bounded. Moreover, the following statement holds
@@
Let $\{\mathbf{w}_{i}^{k}\}$ be the sequence generated by Algorithm 1 and $\sigma\geq r$. Then any accumulation point of $\{\mathbf{w}_{i}^{k}\}$ is a P-stationary point and also a uniquely local minimizer of (3).
Next, we analyze the complexity of Algorithm 1. Since Algorithm 1 is an online method that processes samples in a streaming manner, its computational complexity per sample is given by $\mathcal{O}\left(m(p^{2}+m^{2})\right)$, where $m$ and $p$ are the numbers of dictionary samples and features respectively. This complexity arises from the following three main components. Firstly, updating $\mathbf{y}_{i}$ primarily involves element-wise addition and a hard threshold operation, resulting in a computational complexity of $\mathcal{O}\left(m\right)$. Secondly, updating $\mathbf{x}_{i}$ is the most computationally intensive step. This step involves solving an equation that requires matrix operations such as calculating $D^{\top}D$, adding a scaled identity matrix, and computing the inverse of the resulting matrix. Hence, the overall complexity for updating $\mathbf{x}_{i}$ is $\mathcal{O}\left(m(p^{2}+m^{2})\right)$. Finally, updating $\mathbf{u}_{i}$ only involves simple vector addition, which has a complexity of $\mathcal{O}\left(m\right)$.
3.2 Dictionary Update Based on Support Points
@@
We begin by recalling the definition of the support points proposed in [30].
@@
For a dataset $Z\in\mathbb{R}^{p\times n}$, the support points of $Z$ are defined as
$\{\boldsymbol{\xi}_{i}\}_{i=1}^{m}=\underset{\mathbf{d}_{1},\ldots,\mathbf{d}_{m}}{\arg\min}\frac{2}{mn}\sum_{i=1}^{m}\sum_{l=1}^{n}\|\mathbf{z}_{l}-\mathbf{d}_{i}\|_{2}-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\|\mathbf{d}_{i}-\mathbf{d}_{j}\|_{2}.$ (21)
@@
Assume the dataset $Z\in\mathbb{R}^{p\times n}$ is drawn from $c$ independent subspaces $\{\mathcal{S}_{i}\}_{i=1}^{c}$ with the corresponding dimension $\{d_{i}=\operatorname{dim}(\mathcal{S}_{i})\}_{i=1}^{c}$. Then, it needs at least $\sum_{i=1}^{c}\left(d_{i}+1\right)$ points to adequately capture the distribution of the entire data.
@@
$\min_{\mathbf{d}_{1},\ldots,\mathbf{d}_{m}}\frac{2}{mn}\sum_{i=1}^{m}\sum_{l=1}^{n}\|\mathbf{z}_{l}-\mathbf{d}_{i}\|_{2}-\frac{1}{m^{2}}\left[\sum_{i=1}^{m}\sum_{j=1}^{m}\left(\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}+\frac{2(\mathbf{d}_{i}-\mathbf{d}^{{}^{\prime}}_{j})^{\top}(\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j})}{\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}}\right)\right].$ (22)
@@
$\begin{split}\min_{\mathbf{d}_{1},\ldots,\mathbf{d}_{m}}\frac{2}{mn}\sum_{i=1}^{m}&\sum_{l=1}^{n}\left[\frac{\|\mathbf{z}_{l}-\mathbf{d}_{i}\|_{2}^{2}}{2\|\mathbf{z}_{l}-\mathbf{d}^{{}^{\prime}}_{i}\|_{2}}+\frac{\|\mathbf{z}_{l}-\mathbf{d}^{{}^{\prime}}_{i}\|_{2}}{2}\right]\\ &-\frac{1}{m^{2}}\left[\sum_{i=1}^{m}\sum_{j=1}^{m}\left(\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}+\frac{2(\mathbf{d}_{i}-\mathbf{d}^{{}^{\prime}}_{j})^{\top}(\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j})}{\|\mathbf{d}^{{}^{\prime}}_{i}-\mathbf{d}^{{}^{\prime}}_{j}\|_{2}}\right)\right].\end{split}$ (23)
@@
Let $\mathcal{D}$ be compact and convex, and $\{D^{k}\}$ be the sequence generated by Algorithm 2. If $D_{0}\subseteq\mathcal{D}$, any accumulation point of $\{D^{k}\}$ is a local minimizer of (21).
@@ 3.3 Online Algorithm with Dictionary Update
Algorithm 3 OENSC Algorithm with Support Points
@@ 7                   Update $D$ by Algorithm 2 with $\hat{Z}$;
Algorithm 4 OENSC Algorithm with Dictionary Update (OENSC-S)
@@ Acknowledgment
According to [17], we only need to verify that $g(X)=\lambda_{1}\lVert X\rVert_{0}+\frac{\lambda_{2}}{2}\lVert X\rVert_{F}^{2}$ satisfies the following two conditions: (1) $g(P^{\top}XT)=g(X)$ for any permutation matrix $P$; (2) $g(X)\geq g(X^{B})$ with $X=\left(\begin{matrix}X_{1}&X_{2}\\ X_{3}&X_{4}\end{matrix}\right)$ and $X=\left(\begin{matrix}X_{1}&0\\ 0&X_{4}\end{matrix}\right)$, in which the equality holds if and only if $X=X^{B}$.
@@
$g(P^{\top}XP)$ $=\lambda_{1}\lVert P^{\top}XP\rVert_{0}+\frac{\lambda_{2}}{2}\lVert P^{\top}XP\rVert_{F}^{2}$
$=\lambda_{1}\sum_{i,j}\mathbb{I}\left[(P^{\top}XP)_{ij}\neq 0\right]+\frac{\lambda_{2}}{2}\sum_{i,j}\left[\left(P^{\top}XP\right)_{i,j}\right]^{2}$
$=\lambda_{1}\sum_{i,j}\mathbb{I}\left[P^{\top}(XP)_{ij}\neq 0\right]+\frac{\lambda_{2}}{2}\sum_{i,j}\left[P^{\top}\left(XP\right)_{i,j}\right]^{2}$
$=\lambda_{1}\sum_{i,j}\mathbb{I}\left[P^{\top}(X)_{ij}\neq 0\right]+\frac{\lambda_{2}}{2}\sum_{i,j}\left[P^{\top}\left(X\right)_{i,j}\right]^{2}$
@@
It follows from (9) that the first equation of (7) can be written as
$Y^{*}_{ij}=\left[\operatorname{Prox}_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(Y^{*}+\alpha U^{*}\right)\right]_{ij}=\begin{cases}0,&(Y^{*}+\alpha U^{*})_{ij}\leq\sqrt{2\alpha\lambda_{1}},\\ (Y^{*}+\alpha U^{*})_{ij},&(Y^{*}+\alpha U^{*})_{ij}>\sqrt{2\alpha\lambda_{1}}.\end{cases}$
@@
$0=Y^{*}-\operatorname{Prox}_{\alpha\lambda_{1}\|\cdot\|_{0}}\left(Y^{*}+\alpha U^{*}\right)=\left(\begin{matrix}Y^{*}_{T}\\ Y^{*}_{\bar{T}}\end{matrix}\right)-\left(\begin{matrix}0\\ Y^{*}_{\bar{T}}+\alpha U^{*}_{\bar{T}}\end{matrix}\right)=\left(\begin{matrix}Y^{*}_{T}\\ \alpha U^{*}_{\bar{T}}\end{matrix}\right),$
where $T=\{(i,j):Y^{*}_{ij}\leq\sqrt{2\alpha\lambda_{1}}\}$ and $\bar{T}$ is the complement set of $T$. Thus, $Y^{*}_{ij}=0$ for any $(i,j)\in T$ and $\left|Y^{*}_{ij}\right|>\sqrt{2\alpha\lambda_{1}}$ for any $(i,j)\in\bar{T}$, which implies the assertion (a). Then the assertion (b) holds by $Y^{*}=X^{*}$. This completes the proof.
@@
By the definition of the proximal operator, it follows from the second equation in (7) that
$0\in\alpha\lambda_{1}\partial\lVert Y^{*}\rVert_{0}+Y^{*}-(Y^{*}+\alpha U^{*}).$
This together with $\alpha>0$ immediately yields that $(Y^{*},X^{*},U^{*})$ is a KKT point. Since $h$ is a strongly convex function, [44, Theorem 3.2] leads to $(Y^{*},X^{*})$ is also a uniquely local minimizer. The proof is completed.
@@
Let $\gamma=\lambda_{min}(D^{\top}D)$, it derives
$\mathcal{L}_{\sigma}(\mathbf{y}_{i}^{k+1},\mathbf{x}_{i}^{k+1},\mathbf{u}_{i}^{k+1})-\mathcal{L}_{\sigma}(\mathbf{y}_{i}^{k+1},\mathbf{x}_{i}^{k+1},\mathbf{u}_{i}^{k})$ $\leq-\frac{1}{\sigma}\|\left(D^{\top}D+\lambda_{2}I\right)\left(\mathbf{x}_{i}^{k+1}-\mathbf{x}_{i}^{k}\right)\|_{2}^{2}$
@@
Secondly, we show that any limit points of $\{\mathbf{w}_{i}\}$ is a P-stationary point of problem (3). By taking the limit on both sides of the update formula for the Lagrangian multiplier, it can be reached that $\mathbf{x}_{i}^{*}-\mathbf{y}_{i}^{*}=0$, which demonstrates the feasibility of $\mathbf{w}_{i}^{*}$. Then, the first equation in (7) can be obtained directly from the $\mathbf{x}_{i}$-subproblem. Furthermore, $\mathbf{y}_{i}^{k_{j}}=\operatorname{Prox}_{\frac{\lambda_{1}}{\sigma}\|\cdot\|_{0}}(\mathbf{x}_{i}^{k_{j}}+\mathbf{u}_{i}^{k_{j}}/\sigma)$ can be derived by taking limits on both sides of (16). Thus, $\{\mathbf{w}_{i}^{*}\}$ is a stationary point of problem (3).